### v1

# Strain effects in a directly bonded diamond-on-insulator substrate

Abstract
The direct bonding process of a diamond-on-insulator (DOI) substrate enables monolithic integration of diamond photonic structures for quantum computing by improving photon collection efficiency and entanglement generation rate between emitters. It also addresses key fabrication challenges, such as robustness, bonding strength, and scalability. In this study, we investigate strain effects in DOI substrates following direct bonding. Strain generation is expected near the diamond-SiO2/Si interface due to thermal expansion coefficient mismatch between the bonded materials. Strain-induced lattice distortions are characterized using nitrogen-vacancy (NV) centers in diamond via optically detected magnetic resonance (ODMR) and photoluminescence (PL) mapping. PL mapping reveals interference fringes in unbonded regions, indicating bonding irregularities. Depth-resolved ODMR measurements show a volumetric strain component increase of $\sim$0.45 MHz and a shear component increase of $\sim$0.71 MHz between the top surface and the DOI interface. However, ODMR signal contrast and peak linewidth remain largely unaffected, suggesting no visible deterioration in the optical properties of the emitters. By combining ODMR and PL mapping, this work establishes a robust methodology for assessing bonding quality and strain impact on NV centers, an essential step toward advancing scalable quantum technologies and integrated photonic circuits.
keywords:
Diamond-on-insulator (DOI), direct bonding, nitrogen-vacancy (NV) centers, optically detected magnetic resonance (ODMR), photoluminescence (PL) mapping, strain effects, thermal expansion mismatch
Ioannis Varveris* Gianni D. Aliberti Tianyin Chen Filip A. Sfetcu Diederik J. W. Dekker Alfred Schuurmans Nikolaj K. Nitzsche Salahuddin Nur Ryoichi Ishihara
[…]
{affiliations}
I. Varveris, G. D. Aliberti, F. Sfetcu, A. Schuurmans, N. K. Nitzsche, Prof. R. Ishihara
QuTech, Delft University of Technology, Lorentzweg 1, 2628 CJ, Delft, The Netherlands
Email: i.varveris@tudelft.nl
I. Varveris, G. D. Aliberti, T. Chen, D. J. W. Dekker, N. K. Nitzsche, Dr. S. Nur, Prof. R. Ishihara
Department of Quantum & Computer Engineering, Delft University of Technology, Mekelweg 4, 2628 CD Delft, The Netherlands
1 Introduction
One of the challenges in achieving robust quantum communication and high-fidelity quantum computing has long been the entanglement generation rate between NV centers, which is limited by the relatively low fraction of emitted coherent photons that can be detected. A solution to address this problem is to integrate NV centers into optical cavities with a high ratio of quality factor versus mode volume. This is done to increase the zero-phonon line (ZPL) emission and improve collection efficiency [1].
However, current applications are limited by the challenges of growing large single crystal diamond (SCD) via chemical vapor deposition (CVD), which restricts its use in integrated photonics [2]. To overcome this, researchers are developing thin film diamond integrated with color centers in photonic crystal (PhC) cavities [3]. Two prominent approaches are hybrid and monolithic all-diamond systems [4]. Hybrid systems use non-diamond materials like silicon nitride or gallium phosphide coupled with diamond color centers for easier manufacturing and integration [5, 6, 7]. Nevertheless, they face issues with limited coupling efficiency. Monolithic all-diamond systems, which entail fabricating optical cavities directly on single crystal diamond substrates, have shown success with not only NV [8], but also SiV [9, 10], and GeV centers [8].
Such a monolithic fabrication technique is the direct bonding of diamond on insulator substrates. Hydrophilic (water-attracting) direct bonding, in particular, is preferred over indirect methods using adhesives, due to its high bonding strength, temperature stability, and industrial scalability [11]. Similar to Silicon-On-Insulator (SOI) substrate bonding [12], Diamond-On-Insulator (DOI) bonding adheres to the same operating principles. After the surface of the diamond and the insulator (in this case SiO2) have been activated, they are then brought into contact, in order for them to create a bond, as a result of van der Waals forces, and form covalent bonds [12]. The material is subsequently annealed in order to enhance the quality of the interface [13].
However, despite the effectiveness of direct bonding and subsequent annealing in enhancing interface quality, this fabrication process can introduce unintended strain effects within the diamond structure. It is very likely that the diamond, having undergone the aforementioned chemical processing, will exhibit further stress effects on the surface which is bonded to the insulator substrate. Furthermore, part of the direct bonding process takes place under a temperature of 200°C. During the cooling period, each of the bonded materials contracts at different rates, according to their unique thermal expansion coefficients. This mismatch between the thermal expansion rates between the diamond and the silicon dioxide/silicon can lead to the generation of strain at their interface [14, 15], and potentially further deep into the diamond lattice.
These strain effects can cause deformations to the crystal lattice, which result in both inhomogeneous broadening and degrade the spin dephasing time of the NV centers. Therefore, it is deemed critical to be able to understand and visually map the effects of strain, providing meaningful insight into the quality of the bonded samples.
2 Direct bonding process overview
Hydrophilic direct bonding is a process that involves creating a bond between surfaces treated to exhibit hydrophilic properties. The first step is surface activation, which prepares the surfaces for subsequent bonding by introducing reactive hydroxyl groups (–OH). For the diamond substrate, this involves thorough cleaning using a Piranha solution, a highly oxidative mixture of sulfuric acid (H2SO4) and hydrogen peroxide (H2O2) in a 3:1 ratio. This treatment results in the termination of the diamond surface with hydroxyl groups (C–OH), enhancing its reactivity, by providing essential reactive sites for bonding with other molecules, or facilitating further chemical reactions [13, 16].
The insulator substrate is composed of a layer of SiO2 on Si. This SiO2/Si substrate undergoes O2 plasma treatment, a common method for cleaning and oxidizing silicon surfaces. This process forms silanol groups Si-OH on the substrate’s surface, which promote adhesion, modify the surface energy, and enable further chemical reactions or the attachment of functional molecules [13, 16].
Once the surfaces are activated, the initial bond formation between the diamond and SiO2/Si substrate occurs through van der Waals forces and hydrogen bonding, facilitated by the presence of water molecules. Post-activation, both surfaces become highly hydrophilic, due to the presence of hydroxyl groups (-OH). This hydrophilicity causes the surfaces to attract and adsorb water molecules from the environment, forming a thin, hydrated layer on each surface. Van der Waals forces and hydrogen bonding bring the surfaces close together and stabilize their initial contact. The hydroxyl groups (C–OH and Si-OH) and the adsorbed water molecules create hydrogen bonds between them, that facilitate alignment and bonding of the surfaces. Hydrogen bonds, although individually weak, contribute significantly to the initial adhesion between the surfaces [17].
After the initial bond formation, the bonded DOI substrate undergoes an annealing process. The annealing temperature and duration are critical parameters for achieving optimal bond strength. During annealing, the bond strength is enhanced through a dehydration reaction [17], where water molecules gradually diffuse through the interface and form a strong covalent bond (C–O–Si), expressed by the reaction seen below:
C-OH + HO-Si → C-O-Si + H2O
Table 1 below summarizes the hydrophilic direct bonding process steps that are performed in this research.
Table 1: Process steps for hydrophilic direct bonding of a DOI substrate
Process Step Parameters / Remarks
Piranha treatment Immerse diamond in a solution of 30 ml H2SO4 (95–97%) and 10 ml H2O2 (31%) for 30 minutes at 75°C.
Water rinsing of diamond Rinse with deionized (DI) water for 5–10 minutes. Do not dry with N2 gas.
SiO2 PECVD Deposit SiO2 on Si wafer using PECVD, then dice into $25\times 25$ mm2 chips.
Water rinsing of SiO2/Si substrate Rinse with DI water for 5 minutes and dry with N2 gas.
O2 plasma activation Oxidize using PVA Tepla tool at 1000 W and O2 flow of 400 ml/min for 5 minutes.
Contacting Bring diamond and SiO2/Si into contact at 20°C and 40% humidity.
Annealing Anneal at 200°C for 24 hours with N2 flow.
3 Strain in NV centers
3.1 Causes of strain in directly bonded DOI substrates
[…]
The annealing process, however, is critical for enhancing the bond strength of the two different substrates. This is achieved through dehydration reactions, which further exacerbate the strain effects. During the dehydration of the interface, water molecules diffuse out and covalent bonds are formed between the diamond and the insulator substrate, leading to a slight atomic rearrangement. As a result, this lattice misalignment contributes to the strain effects in the diamond layer[20]. Moreover, the surface activation steps may contribute to strain effects. During the Piranha or the $\mathrm{O_{2}}$ plasma treatment, unwanted surface defects can be introduced, which may alter the surface properties when bonding takes place, leading to additional strain[21].
3.2 Strain physics in NV centers
The Hamiltonian $\hat{H}$ of the ground-state spin triplet state of the NV center in diamond, in the absence of imperfections (no strain) and external fields, can be described by the electron spin-spin interaction, which arises from second-order perturbation theory applied to the spin-orbit interaction [22]. Defining the $z$-axis along the NV-axis results in the following expression for the Hamiltonian:
$\frac{\hat{H}}{h}=DS_{z}^{2}$ (1)
where $h$ is Planck’s constant, S is the spin operator of the NV center, and $D=2.87$ GHz is the zero-field splitting (ZFS) tensor.
3.2.1 Spin-strain interaction
In reality, numerous imperfections exist within the diamond lattice, making the spin-strain interaction non-negligible. According to Udvarhelyi et al (2018) [23], the spin-strain interaction for the ground-state spin triplet can be described as a deformation in the diamond crystal, causing the strain tensor $\varepsilon_{ij}$ to couple to the NV spin through the symmetry-allowed Hamiltonian:
$\hat{H}_{\varepsilon}=\hat{H}_{\varepsilon}^{0}+\hat{H}_{\varepsilon}^{1}+\hat{H}_{\varepsilon}^{2}$ (2)
where each term couples to different spin transitions. $\hat{H}_{\varepsilon}^{0}$ concerns the $|m_{S}=0\rangle$ to $|m_{S}=0\rangle$ transition (shifting or splitting the $|m_{S}=0\rangle$ level), $\hat{H}_{\varepsilon}^{1}$ the $|m_{S}=0\rangle$ to the $|m_{S}=\pm 1\rangle$ transition, and $\hat{H}_{\varepsilon}^{2}$ the $|m_{S}=+1\rangle$ to the $|m_{S}=-1\rangle$ transition.
In the NV reference frame $(x,y,z)$, aligned such that $z$ is the NV-axis and $(x,y)$ spans the perpendicular plane, the general spin-strain Hamiltonian allowed by the $C_{3v}$ symmetry is written as [23]:
$\begin{array}[]{rl}\hat{H}_{\varepsilon}/h=&\Big{\{}[h_{41}(\varepsilon_{xx}+\varepsilon_{yy})+h_{43}\,\varepsilon_{zz}]S_{z}^{2}\\[6.0pt] &+\frac{1}{2}[h_{26}\,\varepsilon_{zx}-\frac{1}{2}h_{25}(\varepsilon_{xx}-\varepsilon_{yy})]\{S_{x},S_{z}\}\\[6.0pt] &+\frac{1}{2}[h_{26}\,\varepsilon_{yz}+h_{25}\,\varepsilon_{xy}]\{S_{y},S_{z}\}\\[6.0pt] &+\frac{1}{2}[h_{16}\varepsilon_{zx}-\frac{1}{2}h_{15}(\varepsilon_{xx}-\varepsilon_{yy})](S_{y}^{2}-S_{x}^{2})\\[6.0pt] &+\frac{1}{2}[h_{16}\,\varepsilon_{yz}+h_{15}\,\varepsilon_{xy}]\{S_{x},S_{y}\}\Big{\}}\\[6.0pt] &\equiv M_{z}S_{z}^{2}+N_{x}\{S_{x},S_{z}\}+N_{y}\{S_{y},S_{z}\}+M_{x}(S_{y}^{2}-S_{x}^{2})+M_{y}\{S_{x},S_{y}\}.\end{array}$ (3)
where $h_{ij}$ are coupling parameters (in units of GHz/strain) that depend on the microscopic electronic structure of the NV center (see Table 2), $M_{x}$, $M_{y}$, and $M_{z}$ are the strain tensor amplitudes (in MHz) and $N_{x}$, $N_{y}$ are non-diagonal strain terms that induce coupling between $|m_{S}=0\rangle$ and the $|m_{S}=\pm 1\rangle$ states, and $\varepsilon_{ij}$ is the strain tensor, with $\mathbf{u}(\mathbf{r})$ being the displacement field:
$\varepsilon_{ij}=\frac{1}{2}\left(\frac{\partial u_{i}}{\partial x_{j}}+\frac{\partial u_{j}}{\partial x_{i}}\right)$ (4)
If the resulting strain-induced mixing between $m_{S}=0$ and $m_{S}=\pm 1$ states is small compared to the zero-field splitting $D$, the off-diagonal components $N_{x}$, $N_{y}$ may then be treated as a small perturbation and be neglected.
Table 2: Spin-strain ($h$) coupling-strength parameters calculated from density functional theory by Udvarhelyi et al (2018) [23].
Parameter Value (MHz/strain)
$h_{43}$ $2300\pm 200$
$h_{41}$ $-6420\pm 90\phantom{0}$
$h_{25}$ $-2600\pm 80\phantom{0}$
$h_{26}$ $-2830\pm 70\phantom{0}$
$h_{15}$ $5700\pm 200$
$h_{16}$ $19660\pm 90\phantom{0}$
3.2.2 Eigenfrequencies including the Zeeman term
In the presence of a homogeneous magnetic field $\mathit{B}$, the NV Hamiltonian acquires an extra term:
$\frac{\hat{H}_{B}}{h}=\gamma\mathit{B}\cdot S,$ (5)
where $\gamma$ is the electron gyromagnetic ratio, $2.8$ MHz/G. After diagonalizing this matrix and including the spin-strain interaction Hamiltonian, the following equation is obtained:
$f_{\pm}=D+M_{z}\pm\sqrt{(\gamma B_{z})^{2}+M_{x}^{2}+M_{y}^{2}}$ (6)
where $f_{\pm}$ is the frequency of each spin transition. From Equation 6, the following relations can be derived:
$\Delta f_{\mathrm{shift}}=\frac{1}{2}(f_{+}+f_{-})-D=\mathit{M}_{z}$ (7)
where $M_{z}$ is the volumetric (axial) strain amplitude, quantifying lattice compression or expansion along the NV center axis, directly affecting the resonance frequency shift observed in ODMR measurements, and:
$\Delta f_{\mathrm{split}}=f_{+}-f_{-}=2\sqrt{(\gamma B_{z})^{2}+M_{x}^{2}+M_{y}^{2}}$ (8)
where $M_{x}$ and $M_{y}$ represent the shear (transverse) strain components, corresponding to lattice distortions perpendicular to the NV axis. These transverse components primarily influence the magnitude of peak splitting in the ODMR spectra.
Therefore, the effects of strain can be directly characterized by ODMR [24], as lattice distortions lead to measurable frequency shifts and splittings of the spin resonance transitions.
4 Experimental Section
The primary objective of this study is to measure the effects of strain in diamonds that have undergone direct bonding with an insulator ($\mathrm{SiO_{2}/Si}$) substrate. To achieve this, NV centers are used due to their excellent properties that make them highly sensitive to external magnetic fields and strain fields. As already described above, strain can cause splitting and shifts in the energy levels of NV centers.
To enable these measurements, a confocal Optically Detected Magnetic Resonance (ODMR) setup is used. With ODMR, these changes in the energy states and spins can be quantified in the form of the presence of a magnetic, electric, or strain field, or temperature variations. The frequency shift $\Delta f_{\mathrm{shift}}$ and frequency splitting $\Delta f_{\mathrm{split}}$ are related to the strain tensor components $\epsilon_{\mathrm{ij}}$, through Equation 3.
Solving Equation 7 for for $M_{z}$ gives:
$M_{z}=\Delta f_{\mathrm{shift}}$ (9)
while solving Equation 8 for $M_{xy}=\sqrt{M_{x}^{2}+M_{y}^{2}}$ results in:
$M_{xy}=\sqrt{\frac{\Delta f_{\mathrm{split}}^{2}}{4}-(\gamma B_{z})^{2}}$ (10)
Therefore, by measuring the strain amplitudes ($M_{x}$, $M_{y}$, $M_{z}$) through ODMR measurements, one can partially reconstruct the strain tensor. Specifically, the volumetric strain component, $M_{z}$, provides a linear combination of the axial strain terms ($\epsilon_{xx}$, $\epsilon_{yy}$, and $\epsilon_{zz}$), weighted according to their orientation with respect to the NV center’s crystallographic axis. In this research, the difference in values for $M_{xy}$, and $M_{z}$ is acquired, after comparing ODMR-obtained spectra at different depths of the diamond’s volume, as illustrated in Figure 1.
Figure 1: Indicative ODMR spectra obtained at the top and at the back surface.
For the experiment, a $10\times 10\,\mathrm{mm^{2}}$ bulk diamond substrate (supplied by EDP), with a thickness of $\sim 300$ µm is investigated. Let this sample be referred to as $\mathrm{DS_{PB}}$. It is a CVD-grown substrate, cut along the {100} crystallographic plane, that contains randomly distributed ensembles of NV centers with a concentration of roughly 15 ppb. This sample underwent the direct bonding process, but due to its large surface area and surface roughness, the result was partial bonding of the DOI substrate, meaning that two regions of interest are present—both bonded and unbonded—on the same sample. Different strain fields may exist in these regions, since only part of its surface formed a strong bond with the insulator substrate underneath it. This bonded region has potentially experienced increased strain effects compared to the unbonded one. By comparing the ODMR peak splitting and frequency shift values at the top surface and at the bottom surface (diamond-$\mathrm{SiO_{2}/Si}$ interface), an overall increase in strain is observed in the interface.
These samples had a few microns etched away from their top surface with $\mathrm{O_{2}}$ plasma, in order to allow focusing on their back surface (diamond-$\mathrm{SiO_{2}/Si}$ interface). This challenge arose due to the objective lens’ small working distance of 400 µm and the 100 µm diameter of the microwave delivery wire, which is placed between the objective and the piezo stage. Although the refractive index of diamond, $n=2.4$, provides some leniency by modifying the effective distance of the objective-to-diamond, further etching of the surface ensures that proper focus on the back surface can be achieved and eliminates any worries of collision of the diamond with the objective during 2D scanning. It should be noted that the following measurements are performed in the presence of an external bias magnetic field $B_{z}=15.7$ G (or 43.9 MHz splitting) along the $z$-axis. A 532 nm green laser is used for optical excitation, at a power of 0.25 mW, and the emitted fluorescence is collected via a single-photon counting module (COUNT-10C-FC), coupled to a fiber.
5 Results and Discussion
Figure 2: Coarse PL maps of the partially-bonded diamond substrate. a) Top surface, b) Back surface.
In Figure 2, differences in the acquired photoluminescence (PL) maps of the top and back surfaces are visible. During the direct bonding process, even with thorough cleaning of the substrates in a cleanroom environment, trapped particles or residues at the diamond-SiO2/Si interface may still remain, hindering proper bonding and inducing localized strain regions. This effect is particularly pronounced in the sample investigated here, owing to its large size of $10\times 10\,\mathrm{mm^{2}}$.
The bonded and unbonded regions can be clearly distinguished, providing a useful tool for evaluating the success and quality of the direct bonding process. Specifically, interference fringes are more apparent in the PL map closer to the bottom surface. This may indicate the presence of particles or impurities that were either not thoroughly rinsed or introduced during one of the direct bonding steps, and subsequently became trapped at the interface. Such particles could hinder proper adhesion and potentially lead to localized unbonded regions. These regions are linked to non-uniform bonding or strain effects that can alter the observed layer thickness and local refractive index. The direct bonding quality has to be improved, especially regarding substrates with larger surfaces. This would enable high-quality, reproducible bonding and more precise strain mapping.
Figure 3: ODMR 2D scans performed in a $60\times 60$ µm2 area, depicting the effects of the volumetric strain amplitude $M_{z}$. a) Top surface, b) Back surface.
Afterward, 2D ODMR maps were obtained, starting with a $60\times 60$ µm2 area at the top surface, and then focusing on the diamond plane close to the diamond-$\mathrm{SiO_{2}/Si}$ interface. Figure 3 represents volumetric strain variation, obtained by calculating the frequency shift $\Delta f_{\mathrm{shift}}$ of every point in the map, relative to the zero field splitting (ZFS) value of $D=$ 2.87 GHz, and solving for $M_{z}$, as seen in Equation 9.
The milder volumetric strain effects at the top surface are considered to be part of the intrinsic strain of the diamond. However, moving closer towards the back surface shows increased strain effects, which are assumed to have been caused by the thermal expansion mismatch that occurs during the direct bonding process.
Figure 4: ODMR 2D scans performed in a $60\times 60$ µm2 area, depicting the effects of the transverse strain amplitude $M_{xy}$. a) Top surface, b) Back surface.
Similarly, Figure 4 demonstrates the spatial variation of shear (transverse) strain, obtained via 2D ODMR mapping by analyzing the peak splitting at each measurement point according to Equation 10. Calculating the transverse strain amplitude ($M_{xy}$) reveals the magnitude of lattice distortions in planes perpendicular to the NV axis (i.e., the $xy$, $yz$, and $xz$ planes). An increase in shear strain is observed near the diamond–SiO2/Si interface, consistent with expectations based on the thermal expansion coefficient mismatch between the bonded materials.
Figure 5: Data from different ODMR graphs, obtained at varying planes of the diamond substrate. The data is averaged out of nine different ODMR scans, each starting from a different region of the surface of the diamond. a) Shows the increase of the $M_{z}$ parameter, acquired via frequency shift compared to the default ZFS value in ODMR measurements, and corresponding to volumetric strain, and b) shows the increase of the $M_{xy}$ parameter, acquired via peak splitting in ODMR measurements, and corresponding to shear strain components.
Next, ODMR measurements were performed at various depths ($z$-axis scans), starting from the top surface of the $\mathrm{DS_{PB}}$ sample and finishing close to the bottom surface (diamond-SiO2/Si interface). These ODMR measurements provide information on multiple quantities, such as the frequency shift, which is related to volumetric strain, the frequency splitting, related to shear strain, ODMR peak contrast, which provides insight into the signal-to-noise ratio (SNR) and peak broadening. These measurements include data from both bonded and seemingly-unbonded regions.
In Figure 5, a gradual increase in both volumetric and shear strain can be seen. Contrary to the 2D ODMR maps discussed earlier, such a measurement allows a direct observation of how the strain effects are increasingly more intense close to the interface. Values of $M_{z}$ range from 0.11 MHz up to 0.56 MHz, whereas values of $M_{xy}$ range from 10.05 MHz to 10.76 MHz.
Figure 6: Data from different ODMR graphs, obtained at varying planes of the diamond substrate. The data is averaged out of nine different ODMR scans, each starting from a different region of the surface of the diamond. a) Shows the decline in average ODMR intensity contrast. b) Shows the decline in the FWHM of the ODMR peaks.
The contrast of the peaks in an ODMR graph can provide important information on the quality of the color centers. It not only regards the signal-to-noise ratio (SNR), but is also affected by effects of decoherence and inhomogeneous broadening, which significantly impact the optical properties of the emitters. These phenomena can be caused by many factors, including the presence of strain fields which are investigated here. Figure 6 demonstrates how contrast is exacerbated close to the diamond-SiO2/Si interface. The initial ODMR intensity contrast at the top surface is approximately 5.32%, going up to an average of 5.51% in the region where the NV centers exhibit the best signal, and finally dropping down to approximately 5.15% close to the interface.
The average Full Width Half Maximum (FWHM) of the peaks in the ODMR spectra follows a similar trend, as it seems to diminish close to the interface. It drops from an average of $6.08$ MHz close to the top surface, down to an average of $5.70$ MHz close to the interface. This is an unexpected result, as one would anticipate that the observed linewidth worsens under increased strain conditions, similar to how contrast is exacerbated.
Regarding the experimental setup, problems occasionally arise when focusing on the back surface of a thick substrate, due to limits set by the working distance of the objective lens, as well as obstruction by the microwave wire. This creates the need for thinning of the diamond layer by O2 plasma etching; a rather time-consuming process. Last but not least, the ODMR measurement process itself is another factor to take into consideration concerning the quality of the measurements. Long scanning durations make the samples more prone to thermal drifts and can introduce experimental parameter fluctuations. The time inefficiency is further exacerbated by the long direct bonding processing time of approximately a week. It is worth noting that future DOI substrates may be comprised of diamonds with a thickness of a few micrometers or even in the nanometer range. In such cases, the aforementioned strain effects and the quality of the emitters are bound to differ and be more localized.
6 Conclusion
Strain characterization was performed on a diamond directly bonded to an insulator substrate (SiO2/Si). The mismatch in thermal expansion coefficients between diamond and SiO2/Si during the direct bonding process induces strain effects, particularly pronounced near the interface. In this study, strain properties were evaluated across regions exhibiting different bonding conditions by conducting depth-resolved measurements using optically detected magnetic resonance (ODMR). The ODMR technique allowed for the correlation of resonance frequency shifts and splittings to volumetric and shear strain amplitudes, respectively. Results revealed increased lattice deformations when approaching the diamond–SiO2/Si interface, with the volumetric strain parameter exhibiting an increase of 0.45 MHz and the shear strain parameter rising by 0.71 MHz. Additionally, photoluminescence (PL) mapping of partially bonded diamond areas identified interference fringes that likely indicate localized strain or imperfect bonding at the interface. In this investigation, only minor changes in ODMR contrast and linewidth were observed, suggesting negligible deterioration in emitter quality. Specifically, contrast deteriorated by approximately 0.36%, while unexpectedly, the linewidth (FWHM) also decreased by 0.38 MHz. In conclusion, we have determined that PL mapping serves as an effective preliminary method for assessing bonding quality, while ODMR provides detailed insights into the magnitude and nature of strain and its impact on NV centers. Nevertheless, further detailed measurements are required to fully understand how strain influences the optical and electronic properties of color centers in directly bonded diamond substrates.
Conflict of Interest The authors declare no conflict of interest.
Data Availability Statement The data that support the findings of this study are available from the corresponding authors upon reasonable request.
Acknowledgements We gratefully acknowledge support from the joint research program “Modular quantum computers” by Fujitsu Limited and Delft University of Technology, co-funded by the Netherlands Enterprise Agency under project number PPS2007.